divide and simplify 3x-18/10 / x-6/14x
Answer by Mr_Mathman(5)

3x-18/10 / x-6/14x = (30x-18)/10 / (14x^2-6)/14x = [14x(30x-18)]/[10(14x^2-6)]
[84x(5x-1)]/[20(7x^2-3)] = [21x(5x-1)]/[5(7x^2-3)]